# Microphone Array Geometry Independent Multi-Talker Distant ASR: NTT System for the DASR Task of the CHiME-8 Challenge

### v1

@@ 3 Speaker diarization module
Figure 2: Schematic diagram of our proposed speaker diarization system. Double-lined arrows represent multi-channel data flow, while single-lined arrows represent single-channel data flow.
This section details our speaker diarization system defined in Eq. (3). Figure 2 is a schematic diagram of the entire diarization system. It consists of three main parts: i) EEND-VC-based pre-diarization, ii) multi-channel speaker counting, and iii) TS-VAD refinement. TS-VAD is a very powerful approach for dealing with the challenging recording conditions of the CHiME challenge [39, 17]. However, it requires initialization, which is usually performed using a simple speaker embedding clustering approach. For improved performance, we propose using a strong EEND-VC system instead. We refer to the EEND-VC pre-diarization as DIA1 and its combination with TS-VAD refinement as DIA2. We also built a third diarization system, DIA3, which is the same as DIA2, but with a stronger TS-VAD system with more training data and longer training epochs.
In the following subsections, we describe the EEND-VC pre-diarization in Section 3.1, the multi-channel speaker counting in Section 3.2, TS-VAD refinement in Section 3.4, and finally describe the details of the settings and training schemes in Section 3.5.
@@ 3.1 EEND-VC-based pre-diarization (DIA1)
Figure 3: Schematic diagram of the EEND-VC applied to the $m$-th channel audio. For simplicity, this shows an example with three distinct speakers (represented by the red, green, and blue waveforms in the bottom visualization). The first chunk $\mathbf{y}_{m,1}$ contains three active speakers in total (red, green, and blue speakers), while the $i$-th chunk $\mathbf{y}_{m,i}$ only contains two active speakers (blue and red speakers).
We use a modified version of the EEND-VC framework [40] as the pre-diarization system, incorporating the following three changes. First, we perform SE using GSS on the local segments found by the EEND segmentation to improve the reliability of the speaker embeddings. Second, we perform EEND-VC per microphone and combine the results with DOVER-Lap, which we found an effective way to exploit the diversity of the distributed microphone recordings. Finally, we decouple speaker counting from EEND-VC, which allows more precise speaker counting.
Figure 3 illustrates an example of the data flow in our EEND-VC-based pre-diarization. The processing of the EEND-VC module consists of several steps, applied separately to each microphone, yielding microphone-dependent diarization results, which consist of estimated utterance boundaries, $\hat{\mathbf{B}}_{m}=\{\hat{\mathbf{B}}_{m,s}\}_{1\leq s\leq\hat{S}}$. First, the speech signal is divided into chunks. Then, we use an NN to estimate the speaker activity for each chunk, which we refer to as EEND-based segmentation. We then perform chunk-wise SE to enhance the speech given to the speaker activities of each detected speaker in each segment, reducing the impact of noise, reverberation, and interfering speakers. Next, we compute speaker embeddings from the enhanced speech signals. Finally, we perform clustering and stitching to assemble the global diarization output together. This process is performed for each channel independently, and then the results are combined using DOVER-Lap as explained in Section 3.3.
@@ 3.1.1 EEND-based local segmentation
First, we divide the microphone signal into chunks of $T_{\mathrm{c}}$ seconds, resulting in a set of $I$ chunks $\{{\mathbf{y}}_{m,i}\in\mathbb{R}^{{T_{\mathrm{c}}^{\mathrm{wav}}}}\}_{1\leq i\leq I}$ for each microphone $m$, where each chunk contains ${T_{\mathrm{c}}^{\mathrm{wav}}}=f_{\mathrm{s}}\cdot T_{\mathrm{c}}$ samples and $f_{\mathrm{s}}$ is the sampling frequency. In our system, we set $T_{\mathrm{c}}=30$ seconds. We then apply EEND separately to each channel of each audio chunk $\mathbf{y}_{m,i}$, yielding local speaker activity posteriors as,
@@
where $\hat{\mathbf{a}}_{m,i,s}\coloneqq[\hat{a}_{m,i,s,t=1},\dots,\hat{a}_{m,i,s,t={T_{\mathrm{c}}^{\mathrm{act}}}}]^{\mathsf{T}}\in[0,1]^{{T_{\mathrm{c}}^{\mathrm{act}}}}$ represents the posterior probabilities of the $s$-th local speaker’s speech activity at each frame index $t$ within the $i$-th chunk; $f_{\mathrm{EEND}}(\cdot)$ is the EEND function implemented with an NN. Note that the length of speaker activity vectors, ${T_{\mathrm{c}}^{\mathrm{act}}}$, is shorter than ${T_{\mathrm{c}}^{\mathrm{wav}}}$ due to the downsampling of EEND. $S_{\mathrm{local}}$ denotes the number of local speakers, a hyper-parameter of EEND that determines the maximum number of speakers present within each chunk.
@@ 3.1.2 Segment-wise speaker embedding computation
We calculate $D$-dimensional speaker embeddings $\{\mathbf{e}_{m,i,s}\in\mathbb{R}^{D}\}_{1\leq s\leq S_{\mathrm{local}}}$ for each local speaker $s$ in each chunk $i$ from the $m$-th microphone given the local speaker activity $\hat{\mathbf{A}}_{m,i}$ obtained from Eq. (6). Most EEND-VC-based approaches, including our previous CHiME-7 system [36] and CHiME-7/8 baseline systems [2, 3], extract speaker embeddings exclusively from non-overlapping intervals where only a single speaker is present as,
@@
and $f_{\mathrm{Emb}}(\cdot)$ is a NN, which maps a speech signal to a speaker embedding vector of dimension $D$, such as ECAPA-TDNN [41]. $\mathbbm{1}(\mathrm{condition})$ denotes an indicator function that returns 1 if the condition is satisfied and 0 otherwise; $\tilde{\mathbf{a}}_{m,i,s}\coloneqq[\tilde{a}_{m,i,s,t=1},\dots,\tilde{a}_{m,i,s,t={T_{\mathrm{c}}^{\mathrm{act}}}}]^{\mathsf{T}}\in\{0,1\}^{{T_{\mathrm{c}}^{\mathrm{act}}}}$ is the frame-wise binary mask to select the non-overlapping frames, $\mathrm{upsample}(\cdot)$ is the linear interpolation used to upsample the length of speaker activity from ${T_{\mathrm{c}}^{\mathrm{act}}}$ to ${T_{\mathrm{c}}^{\mathrm{wav}}}$, $\theta_{\mathrm{emb}}$ is the speaker activity threshold for embedding extraction, and $\odot$ is the Hadamard product. This strategy efficiently captures speaker characteristics by focusing on single-speaker intervals.
However, it becomes challenging to extract speaker embeddings from highly overlapping segments that contain only a few brief single-speaker intervals. To address this, our previous system used a very long chunk size (e.g., 80 seconds), setting the number of local speakers to the maximum anticipated for the use case scenario (four in the case of CHiME-7 [19]). However, when using long chunks, the number of local speakers in a chunk at inference time may exceed the anticipated maximum number of speakers. Besides, using excessively long chunks can degrade EEND performance by increasing the risk of speaker permutation within chunks [36].
In response, we propose a more robust speaker embedding strategy that performs SE before speaker embedding extraction, thereby reducing interference from overlapping speakers. Specifically, we apply a chunk-wise SE to enhance each local speaker’s speech $\hat{\mathbf{x}}_{m,i,s}$ from the audio chunk, using their speaker activity $\hat{\mathbf{a}}_{m,i,s}$ and signals from all microphones $m\in\left\{1,\dots,M\right\}\eqqcolon\mathcal{M}$, $\mathbf{Y}_{i}=\{\mathbf{y}_{m,i}\}_{1\leq m\leq M}$. The speaker embeddings are thus computed as follows,
@@ 3.1.3 Vector clustering
Next, vector clustering is applied to the speaker embeddings to assign each local speaker to a global speaker cluster as follows:
@@
We can use different clustering approaches such as k-means clustering, agglomerative hierarchical clustering [42], and spectral clustering [43]. In this paper, we employ spectral clustering with the Ng-Jordan-Weiss algorithm [44]. We use a radial basis function kernel with $\gamma=1.0$, applied to the cosine similarity between speaker embeddings to compute the affinity matrix. We implement spectral clustering using scikit-learn==1.5.2, introducing a small modification to enforce the constraint that speaker embeddings from local activities within the same chunk should not be clustered into the same cluster [40].
@@ 3.1.4 Stitching and post-processing
The global speaker activity is then converted into a set of utterance boundaries, $\hat{\mathbf{B}}_{m}$ using the post-processing method from [45]. Specifically, we apply a median filter with a kernel of 25 frames to smooth the global speaker activity and subsequently threshold it with $\theta_{\mathrm{eend}}$, which yields the binarized speaker activity $\{\tilde{\mathbf{a}}^{\mathrm{global}}_{m,s}\in\{0,1\}^{{T_{\mathrm{c}}^{\mathrm{act}}}}\}_{1\leq s\leq\hat{S}}$. Each global speaker’s utterance boundaries are then identified by locating frame indices where the binarized speaker activity changes.
@@ 3.2 Multi-channel speaker counting
Our preliminary experiments indicated that speaker counting tends to underestimate the number of speakers in short sessions. To address this issue, we propose to augment the number of speaker embeddings used for counting by (1) considering shorter chunks and (2) exploiting the multiple microphone signals. By augmenting the number of speaker embeddings, we ensure stable counting performance even for short recordings.
@@
Figure 4 and Algorithm 1 detail the procedure of the proposed multi-channel speaker counting scheme. The input of our speaker counting scheme is a set of $M$-channel observed signals $\{\mathbf{y}_{m}\}_{1\leq m\leq M}$, a set of $M$-channel enhanced speech signals divided in chunks $\{\hat{\mathbf{x}}_{m,i,s}\}_{1\leq m\leq M}$ estimated by Eq.(10) and the corresponding local speaker activities $\{\hat{\mathbf{a}}_{m,i,s}\}_{1\leq m\leq M}$ estimated by Eq.(6) for all local speakers and chunks. We first divide these enhanced speech chunks and speaker activities into subchunks of $T_{\mathrm{subc}}$ seconds, producing $\hat{\mathbf{x}}^{\mathrm{subc}}_{m,i,s}\in\mathbb{R}^{T^{\mathrm{wav}}_{\mathrm{subc}}}$ and $\hat{\mathbf{a}}^{\mathrm{subc}}_{m,i,s}\in[0,1]^{T^{\mathrm{act}}_{\mathrm{subc}}}$ for $1\leq i\leq I_{\mathrm{subc}}$ (line 2 in Algorithm 1). $T^{\mathrm{wav}}_{\mathrm{subc}}$ and $T^{\mathrm{act}}_{\mathrm{subc}}$ is the number of samples and frames in each subchunk, and $I_{\mathrm{subc}}$ is the number of subchunks defined as $I_{\mathrm{subc}}=I\times\frac{T_{\mathrm{c}}}{T_{\mathrm{subc}}}$.
We then extract speaker embeddings $\{\mathbf{e}^{\mathrm{subc}}_{m,i^{\prime},s}\in\mathbb{R}^{D}\}_{1\leq s\leq S_{\mathrm{local},i^{\prime}}}$ for each $i^{\prime}$-th subchunk and $m$-th microphone (line 3–4 in Algorithm 1). Note that we exclude the speaker embeddings extracted from subchunks with a total speech duration below a certain threshold $T_{\mathrm{min}}=0.75$ seconds to ensure that only reliable embeddings are used. This results in $S^{\mathrm{subc}}_{i^{\prime}}$ speaker embeddings in each $i^{\prime}$-th subchunk. In all setups, we use $T_{\mathrm{subc}}=15$ second subchunks for speaker counting and $T_{\mathrm{c}}=30$ second chunks for vector clustering, which enables us to double the number of speaker embeddings in the speaker counting step.
@@
Subsequently, we further increase the number of embeddings by exploiting the speaker embeddings obtained from all microphones. There are various methods to utilize multi-channel data for speaker counting. One approach involves performing a single clustering step using speaker embeddings from all microphones. However, channel variations can result in inaccurate estimations of the number of speakers. Another approach is to conduct speaker counting for each microphone individually and then aggregate the results. However, this reduces the number of embeddings for clustering, leading to less reliable results. A compromise consists of performing speaker counting on groups of microphone signals that exhibit minimal variations in their transfer function. This method allows the use of speaker embeddings from relatively similar channels, increasing the number of embeddings while limiting the impact of channel variability. The speaker counts from each microphone group can then be aggregated to achieve the final result.
@@
This approach is implemented as follows. We apply the microphone grouping procedure similar to the one used for speech enhancement [46]. This approach is applied to the multi-microphone signals to identify groups of microphones with similar transfer characteristics. Speaker counting is then performed on each subchunk within these groups (lines 5–6). Specifically, we first calculate a microphone similarity matrix $L=\left[l_{m_{1},m_{2}}\right]\in\left[0,1\right]^{M\times M}$, where each element is the similarity between $m_{1}$-th and $m_{2}$-th microphone observed signals $\mathbf{y}_{m_{1}},\mathbf{y}_{m_{2}}$. The similarity is computed based on the inter-microphone correlations of the first $T_{\mathrm{corr}}$ samples of the signals as follows:
@@
where $\tau$ is a time index and $\mu_{*}$ and $\sigma_{*}$ present the mean and the standard deviation of $\mathbf{y}_{*}$ for $1\leq\tau\leq T_{\mathrm{corr}}$, respectively. We set $T_{\mathrm{corr}}=120$ seconds based on preliminary experimental results. We then apply AHC to the microphones based on the computed similarity matrix, following Ward’s method [42]. This approach determines the optimal number of microphone groups by recursively merging the closest pair of clusters that result in the smallest increase in total within-cluster variance, continuing until the minimum similarity between clusters falls below a predetermined threshold $\theta_{\mathrm{mic}}=0.05$. This process groups microphones into $G$ groups, $\{\mathcal{M}_{g}\}_{1\leq g\leq G}$, of the microphones $\{1,\dots,M\}$.
For each microphone group, $\mathcal{M}_{g}$, we finally estimate the number of speakers $s_{g}$ using the set of speaker embeddings within that group as
@@
Finally, to further improve speaker counting accuracy for each session, we integrate the estimated speaker counts from each microphone group to obtain session-level speaker counts by computing the weighted average of the estimated speaker counts $s_{g}$ for each microphone group $g$, as follows:

### v2

@@ 3 Speaker diarization module
Figure 2: Schematic diagram of proposed speaker diarization system. Double-lined arrows represent multi-channel data flow, while single-lined arrows represent single-channel data flow.
This section details our speaker diarization system defined in Eq. (3). The entire diarization system, as shown in Figure 2, consists of three main parts: i) EEND-VC-based pre-diarization, ii) multi-channel speaker counting, and iii) TS-VAD refinement. TS-VAD is a very powerful approach for dealing with the challenging recording conditions of the CHiME challenge [39, 17], but it requires initialization, which is usually performed using a simple speaker embedding clustering approach. For improved performance, we propose using a strong EEND-VC system instead. We refer to the EEND-VC pre-diarization as DIA1 and its combination with TS-VAD refinement as DIA2. We also built a third diarization system, DIA3, which is the same as DIA2 but with a stronger TS-VAD system featuring more training data and longer training epochs.
In the following subsections, we describe the EEND-VC pre-diarization in 3.1, the multi-channel speaker counting in 3.2, TS-VAD refinement in 3.4, and the details of the settings and training schemes in Section 3.5.
@@ 3.1 EEND-VC-based pre-diarization (DIA1)
Figure 3: Schematic diagram of EEND-VC applied to $m$-th channel audio. For simplicity, this shows an example with three distinct speakers (represented by red, green, and blue waveforms in the bottom visualization). The first chunk $\mathbf{y}_{m,1}$ contains three active speakers in total (red, green, and blue speakers), while the $i$-th chunk $\mathbf{y}_{m,i}$ only contains two active speakers (blue and red).
We use a modified version of the EEND-VC framework [40] as the pre-diarization system, incorporating the following three changes. First, we perform SE using GSS on the local segments found by the EEND segmentation to improve the reliability of the speaker embeddings. Second, we perform EEND-VC for each microphone and combine the results with DOVER-Lap, which we found an effective way to exploit the diversity of the distributed microphone recordings. Finally, we decouple speaker counting from EEND-VC, which allows more precise speaker counting.
Figure 3 shows an example of the data flow in our EEND-VC-based pre-diarization. The processing of the EEND-VC module consists of several steps, applied separately to each microphone, yielding microphone-dependent diarization results that consist of estimated utterance boundaries, $\hat{\mathbf{B}}_{m}=\{\hat{\mathbf{B}}_{m,s}\}_{1\leq s\leq\hat{S}}$. First, the speech signal is divided into chunks and we use an NN to estimate the speaker activity for each chunk, which we refer to as EEND-based segmentation. We then perform chunk-wise SE to enhance the speech given to the speaker activities of each detected speaker in each segment, reducing the impact of noise, reverberation, and interfering speakers. Next, we compute speaker embeddings from the enhanced speech signals. Finally, we perform clustering and stitching to assemble the global diarization output together. This process is performed for each channel independently, and the results are then combined using DOVER-Lap as explained in Section 3.3.
@@ 3.1.1 EEND-based local segmentation
First, we divide the microphone signal into chunks of $T_{\mathrm{c}}$ seconds, resulting in a set of $I$ chunks $\{{\mathbf{y}}_{m,i}\in\mathbb{R}^{{T_{\mathrm{c}}^{\mathrm{wav}}}}\}_{1\leq i\leq I}$ for each microphone $m$, where each chunk contains ${T_{\mathrm{c}}^{\mathrm{wav}}}=f_{\mathrm{s}}\cdot T_{\mathrm{c}}$ samples and $f_{\mathrm{s}}$ is the sampling frequency. In our system, we set $T_{\mathrm{c}}=30$ seconds. We then apply EEND separately to each channel of each audio chunk $\mathbf{y}_{m,i}$, yielding local speaker activity posteriors, as
@@
where $\hat{\mathbf{a}}_{m,i,s}\coloneqq[\hat{a}_{m,i,s,t=1},\dots,\hat{a}_{m,i,s,t={T_{\mathrm{c}}^{\mathrm{act}}}}]^{\mathsf{T}}\in[0,1]^{{T_{\mathrm{c}}^{\mathrm{act}}}}$ represents the posterior probabilities of the $s$-th local speaker’s speech activity at each frame index $t$ within the $i$-th chunk ($f_{\mathrm{EEND}}(\cdot)$ is the EEND function implemented with an NN). Note that the length of speaker activity vectors, ${T_{\mathrm{c}}^{\mathrm{act}}}$, is shorter than ${T_{\mathrm{c}}^{\mathrm{wav}}}$ due to the downsampling of EEND. $S_{\mathrm{local}}$ denotes the number of local speakers, a hyper-parameter of EEND that determines the maximum number of speakers present within each chunk.
@@ 3.1.2 Segment-wise speaker embedding computation
We calculate $D$-dimensional speaker embeddings $\{\mathbf{e}_{m,i,s}\in\mathbb{R}^{D}\}_{1\leq s\leq S_{\mathrm{local}}}$ for each local speaker $s$ in each chunk $i$ from the $m$-th microphone given the local speaker activity $\hat{\mathbf{A}}_{m,i}$ obtained from Eq. (6). Most EEND-VC-based approaches, including our previous CHiME-7 system [36] and CHiME-7/8 baseline systems [2, 3], extract speaker embeddings exclusively from non-overlapping intervals where only a single speaker is present, as
@@
and $f_{\mathrm{Emb}}(\cdot)$ is a NN that maps a speech signal to a speaker embedding vector of dimension $D$, such as ECAPA-TDNN [41]. $\mathbbm{1}(\mathrm{condition})$ denotes an indicator function that returns 1 if the condition is satisfied and 0 otherwise $\tilde{\mathbf{a}}_{m,i,s}\coloneqq[\tilde{a}_{m,i,s,t=1},\dots,\tilde{a}_{m,i,s,t={T_{\mathrm{c}}^{\mathrm{act}}}}]^{\mathsf{T}}\in\{0,1\}^{{T_{\mathrm{c}}^{\mathrm{act}}}}$ is the frame-wise binary mask to select the non-overlapping frames, $\mathrm{upsample}(\cdot)$ is the linear interpolation used to upsample the length of speaker activity from ${T_{\mathrm{c}}^{\mathrm{act}}}$ to ${T_{\mathrm{c}}^{\mathrm{wav}}}$, $\theta_{\mathrm{emb}}$ is the speaker activity threshold for embedding extraction, and $\odot$ is the Hadamard product. This strategy efficiently captures speaker characteristics by focusing on single-speaker intervals.
However, it becomes challenging to extract speaker embeddings from highly overlapping segments that contain only a few brief single-speaker intervals. To address this, our previous system used a very long chunk size (e.g., 80 seconds), setting the number of local speakers to the maximum anticipated for the use case scenario (four in the case of CHiME-7 [19]). However, when using long chunks, the number of local speakers in a chunk at inference time may exceed the anticipated maximum number of speakers. Moreover, using excessively long chunks can degrade EEND performance by increasing the risk of speaker permutation within chunks [36].
We therefore propose a more robust speaker embedding strategy that performs SE before speaker embedding extraction, thereby reducing interference from overlapping speakers. Specifically, we apply a chunk-wise SE to enhance each local speaker’s speech $\hat{\mathbf{x}}_{m,i,s}$ from the audio chunk, using their speaker activity $\hat{\mathbf{a}}_{m,i,s}$ and signals from all microphones $m\in\left\{1,\dots,M\right\}\eqqcolon\mathcal{M}$, $\mathbf{Y}_{i}=\{\mathbf{y}_{m,i}\}_{1\leq m\leq M}$. The speaker embeddings are thus computed, as
@@ 3.1.3 Vector clustering
Next, vector clustering is applied to the speaker embeddings to assign each local speaker to a global speaker cluster as
@@
Various different clustering approaches can be used here, such as k-means clustering, agglomerative hierarchical clustering [42], and spectral clustering [43]. In this paper, we utilize spectral clustering with the Ng-Jordan-Weiss algorithm [44]. We use a radial basis function kernel with $\gamma=1.0$, applied to the cosine similarity between speaker embeddings to compute the affinity matrix. Spectral clustering is implemented using scikit-learn==1.5.2, introducing a small modification to enforce the constraint that speaker embeddings from local activities within the same chunk should not be clustered into the same cluster [40].
@@ 3.1.4 Stitching and post-processing
The global speaker activity is then converted into a set of utterance boundaries $\hat{\mathbf{B}}_{m}$ using the post-processing method from [45]. Specifically, we apply a median filter with a kernel of 25 frames to smooth the global speaker activity and subsequently threshold it with $\theta_{\mathrm{eend}}$, which yields the binarized speaker activity $\{\tilde{\mathbf{a}}^{\mathrm{global}}_{m,s}\in\{0,1\}^{{T_{\mathrm{c}}^{\mathrm{act}}}}\}_{1\leq s\leq\hat{S}}$. Each global speaker’s utterance boundaries are then identified by locating frame indices where the binarized speaker activity changes.
@@ 3.2 Multi-channel speaker counting
Our preliminary experiments indicated that speaker counting tends to underestimate the number of speakers in short sessions. We therefore propose augmenting the number of speaker embeddings used for counting by (1) considering shorter chunks and (2) exploiting the multiple microphone signals. By augmenting the number of speaker embeddings, we ensure stable counting performance even for short recordings.
@@
Figure 4 and Algorithm 1 detail the procedure of the proposed multi-channel speaker counting scheme. The input of our speaker counting scheme is a set of $M$-channel observed signals $\{\mathbf{y}_{m}\}_{1\leq m\leq M}$, a set of $M$-channel enhanced speech signals divided in chunks $\{\hat{\mathbf{x}}_{m,i,s}\}_{1\leq m\leq M}$ estimated by Eq.(10) and the corresponding local speaker activities $\{\hat{\mathbf{a}}_{m,i,s}\}_{1\leq m\leq M}$ estimated by Eq.(6) for all local speakers and chunks. We first divide these enhanced speech chunks and speaker activities into subchunks of $T_{\mathrm{subc}}$ seconds, producing $\hat{\mathbf{x}}^{\mathrm{subc}}_{m,i,s}\in\mathbb{R}^{T^{\mathrm{wav}}_{\mathrm{subc}}}$ and $\hat{\mathbf{a}}^{\mathrm{subc}}_{m,i,s}\in[0,1]^{T^{\mathrm{act}}_{\mathrm{subc}}}$ for $1\leq i\leq I_{\mathrm{subc}}$ (line 2 in Algorithm 1). $T^{\mathrm{wav}}_{\mathrm{subc}}$ and $T^{\mathrm{act}}_{\mathrm{subc}}$ denote the number of samples and frames in each subchunk, respectively, and $I_{\mathrm{subc}}$ is the number of subchunks defined as $I_{\mathrm{subc}}=I\times\frac{T_{\mathrm{c}}}{T_{\mathrm{subc}}}$.
@@
We then extract speaker embeddings $\{\mathbf{e}^{\mathrm{subc}}_{m,i^{\prime},s}\in\mathbb{R}^{D}\}_{1\leq s\leq S_{\mathrm{local},i^{\prime}}}$ for each $i^{\prime}$-th subchunk and $m$-th microphone (lines 3–4 in Algorithm 1). Note that we exclude the speaker embeddings extracted from subchunks with a total speech duration below a certain threshold $T_{\mathrm{min}}=0.75$ seconds to ensure that only reliable embeddings are used. This results in $S^{\mathrm{subc}}_{i^{\prime}}$ speaker embeddings in each $i^{\prime}$-th subchunk. In all setups, we use $T_{\mathrm{subc}}=15$ second subchunks for speaker counting and $T_{\mathrm{c}}=30$ second chunks for vector clustering, which enables us to approximately double the number of speaker embeddings in the speaker counting step. These lengths are determined using the development set in the CHiME-8 challenge. We find that overly small subchunks result in deterioration during the subsequent clustering step, and subchunks with half the length of the chunks perform best across all scenarios. Subsequently, we further increase the number of embeddings by exploiting the speaker embeddings obtained from all microphones. There are various methods to utilize multi-channel data for speaker counting. One approach involves performing a single clustering step using speaker embeddings from all microphones. However, channel variations can result in inaccurate estimations of the number of speakers. Another approach is to conduct speaker counting for each microphone individually and then aggregate the results. However, this reduces the number of embeddings for clustering, leading to less reliable results. A compromise consists of performing speaker counting on groups of microphone signals that exhibit minimal variations in their transfer functions. This method allows the use of speaker embeddings from relatively similar channels, increasing the number of embeddings while limiting the impact of channel variability. The speaker counts from each microphone group can then be aggregated to achieve the final result.
@@
This approach is implemented as follows. First, we apply a microphone grouping procedure similar to the one used for speech enhancement [46]. This approach is applied to the multi-microphone signals to identify groups of microphones with similar transfer characteristics. Speaker counting is then performed on each subchunk within these groups (lines 5–6). Specifically, we first calculate a microphone similarity matrix $L=\left[l_{m_{1},m_{2}}\right]\in\left[0,1\right]^{M\times M}$, where each element is the similarity between $m_{1}$-th and $m_{2}$-th microphone observed signals $\mathbf{y}_{m_{1}},\mathbf{y}_{m_{2}}$. The similarity is computed based on the inter-microphone correlations of the first $T_{\mathrm{corr}}$ samples of the signals, as
@@
where $\tau$ is a time index and $\mu_{*}$ and $\sigma_{*}$ denote the mean and standard deviation of $\mathbf{y}_{*}$ for $1\leq\tau\leq T_{\mathrm{corr}}$, respectively. We set $T_{\mathrm{corr}}=120$ seconds based on preliminary experimental results. We then apply AHC to the microphones based on the computed similarity matrix, following Ward’s method [42]. This approach determines the optimal number of microphone groups by recursively merging the closest pair of clusters that result in the smallest increase in total within-cluster variance, continuing until the minimum similarity between clusters falls below a predetermined threshold $\theta_{\mathrm{mic}}=0.05$. This process groups microphones into $G$ groups, $\{\mathcal{M}_{g}\}_{1\leq g\leq G}$, of the microphones $\{1,\dots,M\}$.
For each microphone group $\mathcal{M}_{g}$, we estimate the number of speakers $s_{g}$ using the set of speaker embeddings within that group, as
@@
Finally, to further improve the speaker counting accuracy for each session, we integrate the estimated speaker counts from each microphone group to obtain session-level speaker counts by computing the weighted average of the estimated speaker counts $s_{g}$ for each microphone group $g$, as